# The Gauss map of a projective foliation

### v2

@@ 3.7.1. Proof of proposition 4
We will apply theorem 3. The idea is to use that $Sing_{2}(\mathcal{F})=\bigcup_{1\leq i<j\leq k}V_{ij}$, where $V_{ij}=(f_{i}=f_{j}=0)$ (cf. [4]). This is true by the assumption that $\omega_{\mathcal{F}}$ has no zero divisors.
@@
Now, since $\omega_{\mathcal{F}}$ has no zero divisor, we have $deg(\mathcal{F})=\sum_{i=1}^{k}p_{i}-2$ (see [27]) and
@@
where $deg(f_{j})=p_{j}$, $1\leq i\leq k$. If $\omega$ is generic, then the hypersurfaces $(f_{i}=0)$ and $(f_{j}=0)$ are transverse, for all $i<j$, so that we can apply proposition 4. In particular, $\mathcal{G}_{\mathcal{F}}$ is birational if, and only if,
@@
Therefore, in $\Delta$, $h_{k}$ is strictly increasing with respect to $x_{j}$, $\forall j$. In particular, in $\Delta$ we have
@@
It remains to prove that if $k\leq 3$ then $k=2$ and $(x_{1},x_{2})=(1,2)$. In fact, we have $h_{3}(1,1,1)=g(3)=0$ and $h_{3}(1,1,2)=2$, so that $h_{3}(x_{1},x_{2},x_{3})\neq 1$ $\forall(x_{1},x_{2},x_{3})\in\Delta$. Therefore $k=2$ as we wished. Finally, if $x_{1}>1$ or $x_{2}>2$ then $h_{2}(x_{1},x_{2})>h_{2}(1,2)=1$, so that $(x_{1},x_{2})=(1,2)$ (see example 2).∎
@@ 4. Exceptional components
In homogeneous coordinates $(x,y,z,t)$ the generic exceptional foliation in $\mathcal{E}(2,3;3)$ is defined by the form
@@
We would like to observe that all these three hypersurfaces are $\mathcal{E}$-invariant. In fact, they are branches of levels of the first integral $F=f^{2}/g^{3}$, where $f(x,y,z,t)=zt^{2}-xyt+x^{3}/3$ and $g(x,y,z,t)=yt-x^{2}/2$. Since $\mathcal{E}(2,3;n)$ is the set of linear pull-back foliations of $\mathcal{E}(2,3;3)$ by a linear map $L\colon\mathbb{P}^{n}\dasharrow\mathbb{P}^{3}$, by remark 2.2, if $\mathcal{F}\in\mathcal{E}(2,3;n)$ and $n\geq 4$ then $\mathcal{G}_{\mathcal{F}}$ cannot be dominant.
4.2. Gauss map for the foliations in the boundary of the exceptional component $\mathcal{E}(2,3;3)$
@@ 4.2.1. The boundary of $\mathcal{E}(2,3;3)$
Here we will use this fact to determine some representative elements of the boundary of $\mathcal{E}(2,3;3)$. A well-known result says that every 1-parameter subgroup of $SL_{4}(\mathbb{C})$ is diagonalizable. This type of subgroup is conjugated to one given as below:
@@
The 1-parameter subgroup acts on the variables as:
@@
In order to see how $\lambda(n_{1},n_{2},n_{3})$ acts on $\mathcal{E}$ we set $\omega=\omega_{1}+\omega_{2}+\omega_{3}$, where
@@
    $\omega_{1}^{\prime}+\omega_{2}^{\prime}$ $=-3c^{2}e\,da+3bce\,db+(ace-2b^{2}e)\,dc+(2ac^{2}-b^{2}c)\,de$
@@
    $\omega_{1}^{\prime}+\omega_{3}^{\prime}=c(ab-3ec)\,da-2a^{2}cdb+a(ec+ab)\,dc+2ac^{2}\,de$
@@
    $\omega_{2}^{\prime}+\omega_{3}^{\prime}=(-abc)\,da+c(2a^{2}-3be)\,db+b(-a^{2}+2be)\,dc+(b^{2}c)\,de$
@@
As we saw before, they have birational Gauss map but they are not in the logarithmic components where the generic foliation has birational Gauss map (see corollary 2.5).
@@
We have that $\omega_{1}+\omega_{3}$ is a self-dual foliation. The foliation $\omega_{1}+\omega_{2}$ is projectively equivalent to $\omega^{\prime}_{2}+\omega^{\prime}_{3}$; and the foliation $\omega_{2}+\omega_{3}$ is projectively equivalent to $\omega^{\prime}_{1}+\omega^{\prime}_{2}$
@@
$\omega^{\prime}_{1}+\omega^{\prime}_{2}=$ $ce(2ac-b^{2})\Big{(}2\frac{dc}{c}+\frac{de}{e}-\frac{3}{2}\frac{d(2ac-b^{2})}{2ac-b^{2}}\Big{)},$
$\omega^{\prime}_{2}+\omega^{\prime}_{3}=$ $bc(a^{2}-2be)\Big{(}-2\frac{db}{b}+\frac{dc}{c}+\frac{1}{2}\frac{d(a^{2}-2be)}{a^{2}-2be}\Big{)},$
$\omega^{\prime}_{1}+\omega^{\prime}_{3}=$ $ac(ce-ab)\Big{(}-3\frac{da}{a}-\frac{dc}{c}+2\frac{d(ce-ab)}{ce-ab}\Big{)}.$
@@
$\lim_{s\to\infty}\beta(s)\omega=yt(-3x+2y)dx+xt(x-y+t)dy+y(x^{2}-xy-yt)dt,$
@@
Here we generalize the exceptional components $\mathcal{E}(2,3;3)\subset\mathbb{F}\text{ol}(2;3)$ and $\mathcal{E}(3,4;4)\subset\mathbb{F}\text{ol}(3;4)$. A typical foliation on this component has a rational first integral of the form $f^{n-1}/g^{n}$, where $f$ and $g$ are irreducible and $deg(f)=n$ and $deg(g)=n-1$. It will be generated by a Lie algebra of dimension $n$ of linear vector fields on $\mathbb{C}^{n+1}$
Lemma 4.1.
@@
We assert that $\mathcal{G}:=\left<R,S,N,N^{2},...,N^{n}\right>_{\mathbb{C}}$ is a Lie algebra and $\mathcal{G}_{k}:=\left<R,S,N,N^{2},...,N^{k}\right>_{\mathbb{C}}$ is a Lie sub-algebra of $\mathcal{G}$, for all $k\geq 1$.
@@
Proof. A direct computation shows that $[S,N]=-N$ and that $[R,L]=0$ for all linear vector field $L\in\mathcal{L}(n+1,\mathbb{C})$. On the other hand, from $SN-NS=[S,N]=-N$ we get $SN=NS-N$ and $NS=SN+N$. From this we get
$[S,N^{2}]=SN^{2}-N^{2}S=(NS-N)N-N(SN+N)=-2N^{2}\,.$
@@
$[S,N^{k+1}]=SN^{k+1}-N^{k+1}S=(SN)N^{k}-N^{k}(NS)=(NS-N)N^{k}-N^{k}(SN+N)=$
$=NSN^{k}-N^{k}SN-2N^{k+1}=N(SN^{k-1}-N^{k-1}S)N-2N^{k-1}=$
$=N[S,N^{k-1}]N-2N^{k+1}$
Suppose by induction on $j\geq 1$ that $[S,N^{j}]=-jN^{j}$, $\forall 1\leq j\leq k$, where $k\geq 2$. Then the above computation implies that
$[S,N^{k-1}]=-(k-1)N^{k-1}\,\implies\,[S,N^{k+1}]=-(k+1)N^{k+1}\,.$
This proves that $\mathcal{G}$ is a Lie algebra and that $\mathcal{G}_{k}$ is a Lie sub-algebra of $\mathcal{G}$, $\forall k\geq 1$. ∎A consequence of lemma 4.1 is that the 1-form
@@
is integrable: $\omega\wedge d\omega=0$. Moreover, $\omega$ defines a codimension one foliation $\mathcal{F}_{\omega}$, of degree $n-1$ of $\mathbb{P}^{n}$, because their coefficients are homogeneous of degree $n$. The foliation $\mathcal{F}_{\omega}$ is generated on homogenous coordinates by the Lie algebra $\mathcal{G}_{n-2}$.
@@
The Gauss map of the foliation $\mathcal{F}_{\omega}$ is birrational. Moreover it has a rational first integral of the form $f^{n-1}/g^{n}$, where $deg(f)=n$ and $deg(g)=n-1$.
@@
where $\mu=\frac{\partial}{\partial z_{1}}\wedge...\wedge\frac{\partial}{\partial z_{n}}$. Note that $f$ is homogeneous of degree $n$ and $g$ of degree $n-1$. We assert that $S(f)=n\,f$, $S(g)=(n-1)\,g$ and $N(f)=N(g)=0$. We will use that
@@
Relation (7) implies that $L_{S}(\mu)=-tr(S)\mu$, $tr(S)=1+2+...+n$, so that:
@@
On the other hand, by lemma 4.1 we have:
@@
$S(f)=n\,f$.
In the case of $g$, we have $L_{S}(\rho)=S(g)\mu-tr(S)g\mu$ and from lemma 4.1:
@@
$f^{n-1}/g^{n}$ is a first integral of $S$. Since $[N,\frac{\partial}{\partial z_{0}}]=[N,\frac{\partial}{\partial z_{n}}]=0$, from (7) we get $N(f)=N(g)=0$, so that $N(f^{n-1}/g^{n})=0$. Finally, since $deg(f^{n-1})=deg(g^{n})=n(n-1)$ we get $R(f^{n-1}/g^{n})=0$ and $f^{n-1}/g^{n}$ is a first integral of all the generators of $\mathcal{G}_{n-2}$, and so of the form $\omega$.
Remark 4.1.
@@
Using ()(*)( ∗ ) and ()(**)( ∗ ∗ ) we obtain that $f(0,z_{1},...,z_{n})=z_{1}^{n}$ and $g(0,z_{1},...,z_{n})=z_{1}^{n-1}$, so that $(z_{0}=0)$ is contained in the leaf $f^{n-1}-g^{n}=0$.
@@
On the other hand, from ()(*)( ∗ ) and ()(**)( ∗ ∗ ) we get
@@
Proof. The idea is to prove that $\mathcal{F}_{\omega}$ satisfies the hypothesis of corollary 2.4 of theorem 3. Let us compute the irreducible components of $Sing_{2}(\mathcal{F}_{\omega})$. First of all, from lemma 4.2 we have $\omega|_{(z_{0}=1)}=\alpha^{-1}[(n-1)g\,df-nf\,dg]$, so that
@@
$Sing(\mathcal{F}_{\omega})\cap(z_{0}=1)=(f=g=0)\cap(z_{0}=1)$, because $df\wedge dg\not\equiv 0$. Moreover, $Sing(\mathcal{F}_{\omega})\cap(z_{0}=1)$ is a union of (singular) orbits of the action of $\mathcal{G}_{n-2}$. In fact, it is composed of just one orbit: the orbit of the point $p_{o}:=(1,0,...,0)\in(z_{0}=1)$, because
$Sing(\mathcal{F}_{o})\cap(z_{0}=1)\cap(z_{1}=z_{2}=...=z_{n-2}=0)=\{p_{o}\}\,,$
as the reader can check. In particular, $Sing(\mathcal{F}_{\omega})\cap(z_{0}=1)$ is a connected smooth sub-variety of codimension two of $(z_{0}=1)\simeq\mathbb{C}^{n}$. We denote it by $V_{1}$. Since $(f=0)$ cuts transversely $(g=0)$ along $(z_{0}=1)$, the transversal type of $\mathcal{F}_{\omega}$ along $V_{1}$ has linear part $(n-1)u\,dv-nv\,du$, so that $\mu(\mathcal{F}_{\omega},V_{1})=1$.
The varieties $(f=0)$ and $(g=0)$ cut $(z_{0}=0)$ along $V_{2}:=(z_{0}=z_{1}=0)$, but they are tangent along $V_{2}$ in order $n-1$. This implies that, as a scheme, the variety $(f=g=0)$ is $V_{1}\oplus V_{2}^{n}$. In particular, since $deg(f)=n$ and $deg(g)=n-1$, we have $deg(V_{1})=n(n-1)-n=n^{2}-2n$.
Now, from ()($\star$)( ⋆ ) we have $Sing(\mathcal{F}_{\omega})\cap(z_{0}=0)=(z_{1}.\,F=0)$, where
@@
$Sing(\mathcal{F}_{\omega})\cap(z_{0}=0)=V_{2}\cup V_{3}$, where $V_{3}=(F=0)$ and $deg(V_{3})=n-1$.
Finally, since $deg(V_{1})+deg(V_{2})+deg(V_{3})=n^{2}-n=(n-1)^{2}+(n-1)=deg(\mathcal{F}_{\omega})+deg^{2}(\mathcal{F}_{\omega})$ the Gauss map of $\mathcal{F}_{\omega}$ is birational, by corollary 2.4.∎
For instance, in the case $n=5$ we have:
@@
The singular set $Sing(\mathcal{F}_{\omega})$ has three irreducible components, say $V_{1}$, $V_{2}$ and $V_{3}$, where $V_{3}\cap(z_{0}=0)$ is $(f|_{z_{0}=1}=g|_{z_{0}=1}=0)$. The closure of this variety is irreducible of degree $15$. Finally, $Sing(\mathcal{F}_{\omega})\cap(z_{0}=0)=V_{1}\cup V_{2}$, where:
@@
the first one of degree $1$ and the second of degree $4$.

### v3

@@ 3.7.1. Proof of Proposition 4
We will apply Theorem 3. The idea is to use that $Sing_{2}(\mathcal{F})=\bigcup_{1\leq i<j\leq k}V_{ij}$, where $V_{ij}=(f_{i}=f_{j}=0)$ (see [4]). This is true by the assumption that $\omega_{\mathcal{F}}$ has no zero divisors.
@@
Now, since $\omega_{\mathcal{F}}$ has no zero divisor, we have $deg(\mathcal{F})=\sum_{i=1}^{k}p_{i}-2$ (see example 2.5 of [28]) and
@@ 3.7.2. Proof of corollary 6
where $deg(f_{j})=p_{j}$, $1\leq i\leq k$. If $\omega$ is generic, then the hypersurfaces $(f_{i}=0)$ and $(f_{j}=0)$ are transverse, for all $i<j$, so that we can apply the proposition 4. In particular, $\mathcal{G}_{\mathcal{F}}$ is birational if, and only if,
@@
Therefore, in $\Delta$, $h_{k}$ is strictly increasing with respect to $x_{j}$, for all $j$. In particular, in $\Delta$ we have
@@
It remains to prove that if $k\leq 3$ then $k=2$ and $(x_{1},x_{2})=(1,2)$. In fact, we have $h_{3}(1,1,1)=g(3)=0$ and $h_{3}(1,1,2)=2$, so that $h_{3}(x_{1},x_{2},x_{3})\neq 1$, for all $(x_{1},x_{2},x_{3})\in\Delta$. Therefore $k=2$ as we wished. Finally, if $x_{1}>1$ or $x_{2}>2$ then $h_{2}(x_{1},x_{2})>h_{2}(1,2)=1$, so that $(x_{1},x_{2})=(1,2)$ (see example 2).∎
@@ 4. Exceptional components
In homogeneous coordinates $(x,y,z,t)$, the generic exceptional foliation in $\mathcal{E}(2,3;3)$ is defined by the form
@@
We would like to observe that all these three hypersurfaces are $\mathcal{E}$-invariant. In fact, they are branches of levels of the first integral $F=f^{2}/g^{3}$, where $f(x,y,z,t)=zt^{2}-xyt+x^{3}/3$ and $g(x,y,z,t)=yt-x^{2}/2$.
Since $\mathcal{E}(2,3;n)$ is the set of linear pull-back foliations of $\mathcal{E}(2,3;3)$ by a linear map $L\colon\mathbb{P}^{n}\dasharrow\mathbb{P}^{3}$, by remark 2, if $\mathcal{F}\in\mathcal{E}(2,3;n)$ and $n\geq 4$ then $\mathcal{G}_{\mathcal{F}}$ cannot be dominant.
4.2. Gauss map for foliations in the boundary of the exceptional component $\mathcal{E}(2,3;3)$
@@ 4.2.1. The boundary of $\mathcal{E}(2,3;3)$
Here we will use this fact to determine some representative elements of the boundary of $\mathcal{E}(2,3;3)$. First, we must recall that a 1-parameter subgroup of $SL_{4}(\mathbb{C})$ is an algebraic morphism from $\mathbb{C}^{*}$ to the group and, since $\mathbb{C}^{*}$ is abelian, all such subgroups are diagonalizable. Then, a 1-parameter subgroup of $SL_{4}(\mathbb{C})$ is conjugated to one given as below:
@@
This 1-parameter subgroup acts on the variables as:
@@
In order to see how $\lambda_{(n_{1},n_{2},n_{3})}$ acts on $\mathcal{E}$ we set $\omega=\omega_{1}+\omega_{2}+\omega_{3}$, where
@@
    $(\omega_{1}+\omega_{2})^{\prime}$ $=-3c^{2}e\,da+3bce\,db+(ace-2b^{2}e)\,dc+(2ac^{2}-b^{2}c)\,de$
@@
    $(\omega_{1}+\omega_{3})^{\prime}=c(ab-3ec)\,da-2a^{2}cdb+a(ec+ab)\,dc+2ac^{2}\,de$
@@
    $(\omega_{2}+\omega_{3})^{\prime}=(-abc)\,da+c(2a^{2}-3be)\,db+b(-a^{2}+2be)\,dc+(b^{2}c)\,de$
@@
As we saw before, they have birational Gauss map but they are not in the logarithmic components where the generic foliation has birational Gauss map (see corollary 6). With this, we can also see that the divisors contracted by the three Gauss maps of the three foliations are invariant to the respective foliations.
@@
We have that $\omega_{1}+\omega_{3}$ is a self-dual foliation. The foliation $\omega_{1}+\omega_{2}$ is projectively equivalent to $(\omega_{2}+\omega_{3})^{\prime}$; and the foliation $\omega_{2}+\omega_{3}$ is projectively equivalent to $(\omega_{1}+\omega_{2})^{\prime}$ .
@@
$(\omega_{1}+\omega_{2})^{\prime}=$ $ce(2ac-b^{2})\Big(2\frac{dc}{c}+\frac{de}{e}-\frac{3}{2}\frac{d(2ac-b^{2})}{2ac-b^{2}}\Big),$
$(\omega_{2}+\omega_{3})^{\prime}=$ $bc(a^{2}-2be)\Big(-2\frac{db}{b}+\frac{dc}{c}+\frac{1}{2}\frac{d(a^{2}-2be)}{a^{2}-2be}\Big),$
$(\omega_{1}+\omega_{3})^{\prime}=$ $ac(ce-ab)\Big(-3\frac{da}{a}-\frac{dc}{c}+2\frac{d(ce-ab)}{ce-ab}\Big).$
@@
$\lim_{s\to\infty}\frac{\beta(s)\cdot\omega}{s^{4}}=yt(-3x+2y)dx+t(x^{2}-xy+yt)dy+y(2x^{2}-xy-yt)dt,$
@@
Here we generalize the exceptional components $\mathcal{E}(2,3;3)\subset\mathbb{F}\text{ol}(2;3)$ and $\mathcal{E}(3,4;4)\subset\mathbb{F}\text{ol}(3;4)$. A typical foliation on this component has a rational first integral of the form $f^{n-1}/g^{n}$, where $f$ and $g$ are irreducible and $deg(f)=n$ and deg(g)=n1deg(g)=n-1. It will be generated by a Lie algebra of dimension $n$ of linear vector fields on $\mathbb{C}^{n+1}$
Lemma 1.
@@
We assert that $\mathcal{G}:=\left<R,S,N,N^{2},...,N^{n}\right$$_{\mathbb{C}}$ is a Lie algebra and $\mathcal{G}_{k}:=\left<R,S,N,N^{2},...,N^{k}\right$$_{\mathbb{C}}$ is a Lie sub-algebra of $\mathcal{G}$, for all $k\geq 1$.
@@
Proof. A direct computation shows that [S,N]=N[S,N]=-N and that $[R,L]=0$ for all linear vector field $L\in\mathcal{L}(n+1,\mathbb{C})$. On the other hand, from SNNS=[S,N]=NSN-NS=[S,N]=-N we get SN=NSNSN=NS-N and $NS=SN+N$. From this we get
$S,N^{2}]=SN^{2}-N^{2}S=(NS-N)N-N(SN+N)=-2N^{2}\,.$
@@
$S,N^{k+1}]=SN^{k+1}-N^{k+1}S=(SN)N^{k}-N^{k}(NS)=(NS-N)N^{k}-N^{k}(SN+N)=$
$NSN^{k}-N^{k}SN-2N^{k+1}=N(SN^{k-1}-N^{k-1}S)N-2N^{k-1}=$
$N[S,N^{k-1}]N-2N^{k+1}$
Suppose by induction on $j\geq 1$ that $S,N^{j}]=-jN^{j}$, $\forall 1\leq j\leq k$, where $k\geq 2$. Then the above computation implies that
$S,N^{k-1}]=-(k-1)N^{k-1}\,\implies\,[S,N^{k+1}]=-(k+1)N^{k+1}\,.$
This proves that $\mathcal{G}$ is a Lie algebra and that $\mathcal{G}_{k}$ is a Lie sub-algebra of $\mathcal{G}$, $\forall k\geq 1$. ∎
A consequence of lemma 1 is that the 1-form
@@
is integrable: $\omega\wedge d\omega=0$. Moreover, $\omega$ defines a codimension one foliation $\mathcal{F}_{\omega}$, of degree n1n-1 of $\mathbb{P}^{n}$, because their coefficients are homogeneous of degree $n$. The foliation $\mathcal{F}_{\omega}$ is generated on homogenous coordinates by the Lie algebra $\mathcal{G}_{n-2}$.
@@
The Gauss map of the foliation $\mathcal{F}_{\omega}$ is birrational. Moreover it has a rational first integral of the form $f^{n-1}/g^{n}$, where $deg(f)=n$ and deg(g)=n1deg(g)=n-1.
@@
where $\mu=\frac{\partial}{\partial z_{1}}\wedge...\wedge\frac{\partial}{\partial z_{n}}$. Note that $f$ is homogeneous of degree $n$ and $g$ of degree n1n-1. We assert that S(f)=nfS(f)=n\,f, S(g)=(n1)gS(g)=(n-1)\,g and $N(f)=N(g)=0$. We will use that
@@
Relation (7) implies that $L_{S}(\mu)=-tr(S)\mu$, tr(S)=1+2++ntr(S)=1+2+...+n, so that:
@@
On the other hand, by lemma 1 we have:
@@
S(f)=nfS(f)=n\,f.
In the case of $g$, we have $L_{S}(\rho)=S(g)\mu-tr(S)g\mu$ and from lemma 1:
@@
$f^{n-1}/g^{n}$ is a first integral of $S$. Since $N,\frac{\partial}{\partial z_{0}}]=[N,\frac{\partial}{\partial z_{n}}]=0$, from (7) we get $N(f)=N(g)=0$, so that $N(f^{n-1}/g^{n})=0$. Finally, since $deg(f^{n-1})=deg(g^{n})=n(n-1)$ we get $R(f^{n-1}/g^{n})=0$ and $f^{n-1}/g^{n}$ is a first integral of all the generators of $\mathcal{G}_{n-2}$, and so of the form $\omega$.
Remark 14.
@@
Using ()(*) and ()(**) we obtain that $f(0,z_{1},...,z_{n})=z_{1}^{n}$ and $g(0,z_{1},...,z_{n})=z_{1}^{n-1}$, so that $(z_{0}=0)$ is contained in the leaf $f^{n-1}-g^{n}=0$.
@@
On the other hand, from ()(*) and ()(**) we get
@@
Proof. The idea is to prove that $\mathcal{F}_{\omega}$ satisfies the hypothesis of Corollary 5 of Theorem 3. Let us compute the irreducible components of $Sing_{2}(\mathcal{F}_{\omega})$. First of all, from Lemma 2 we have $\omega|_{(z_{0}=1)}=\alpha^{-1}[(n-1)g\,df-nf\,dg]$, so that
@@
$Sing(\mathcal{F}_{\omega})\cap(z_{0}=1)=(f=g=0)\cap(z_{0}=1)$, because $df\wedge dg$ is not zero at $z_{0}=1$. Moreover, $Sing(\mathcal{F}_{\omega})\cap(z_{0}=1)$ is a union of (singular) orbits of the action of $\mathcal{G}_{n-2}$. In fact, it is composed of just one orbit: the orbit of the point $p_{o}:=(1,0,...,0)\in(z_{0}=1)$, because
$Sing(\mathcal{F}_{\omega})\cap(z_{0}=1)\cap(z_{1}=z_{2}=...=z_{n-2}=0)=\{p_{o}\}\,$,
as the reader can check. In particular, $Sing(\mathcal{F}_{\omega})\cap(z_{0}=1)$ is a connected smooth sub-variety of codimension two of $(z_{0}=1)\simeq\mathbb{C}^{n}$. We denote it by $V_{1}$. Since $(f=0)$ cuts transversely $(g=0)$ along $(z_{0}=1)$, the transversal type of $\mathcal{F}_{\omega}$ along $V_{1}$ has linear part (n1)udvnvdu(n-1)u\,dv-nv\,du, so that $\mu(\mathcal{F}_{\omega},V_{1})=1$.
The varieties $(f=0)$ and $(g=0)$ cut $(z_{0}=0)$ along $V_{2}:=(z_{0}=z_{1}=0)$, but they are tangent along $V_{2}$ in order n1n-1. This implies that, as a scheme, $(f=g=0)$ has $V_{2}$ appearing with multiplicity $n$, reflecting the higher-order tangency, while $V_{1}$ appears with multiplicity one. In particular, since $deg(f)=n$ and deg(g)=n1deg(g)=n-1, we have $deg(V_{1})=n(n-1)-n=n^{2}-2n$.
Now, from $(\star)$ we have $Sing(\mathcal{F}_{\omega})\cap(z_{0}=0)=(z_{1}.\,F=0)$, where
@@
$Sing(\mathcal{F}_{\omega})\cap(z_{0}=0)=V_{2}\cup V_{3}$, where $V_{3}=(z_{0}=F=0)$ and $deg(V_{3})=n-1$.
Finally, since $deg(V_{1})+deg(V_{2})+deg(V_{3})=n^{2}-n=(n-1)^{2}+(n-1)=deg(\mathcal{F}_{\omega})+deg^{2}(\mathcal{F}_{\omega})$ the Gauss map of $\mathcal{F}_{\omega}$ is birational, by corollary 5.∎
For example, in the case $n=5$ we have:
@@
The singular set $Sing(\mathcal{F}_{\omega})$ has three irreducible components, say $V_{1}$, $V_{2}$ and $V_{3}$, where $V_{3}\cap(z_{0}=1)$ is $(f|_{z_{0}=1}=g|_{z_{0}=1}=0)$. The closure of this variety is irreducible of degree $15$. Finally, $Sing(\mathcal{F}_{\omega})\cap(z_{0}=0)=V_{1}\cup V_{2}$, where:
@@
the first of degree $1$ and the second of degree $4$.